OpenStudy (anonymous):

Why lim x->0 2*tan(x^2)/x^2=2?

OpenStudy (ash2326):

The limit is incorrect, how did you find this?

OpenStudy (anonymous):

@Mertsj what do you think?

OpenStudy (mertsj):

The limit is 2

OpenStudy (anonymous):

explain, please

OpenStudy (anonymous):

it says nothing, we must come up by hand

OpenStudy (mertsj):

Click the show steps button

OpenStudy (mertsj):

It shows every step.

OpenStudy (anonymous):

I think we can break it down as lim 2 sin x^2 /cos x^2 x^2\[= 2\lim \frac{ \sin (x^2) }{ x^2 }\frac{ 1 }{ \cos (x^2) } =2\] what do you think?

OpenStudy (anonymous):

of course x ---> 0

OpenStudy (mertsj):

I agree;.

OpenStudy (anonymous):

bingo. sorry Mertsji, I have to do every thing by hand, so I am better not use the math tool

OpenStudy (mertsj):

it is a good practice.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Just another approach.

OpenStudy (anonymous):

you are right!!!

OpenStudy (anonymous):

lol...thanks Hoa! Keep up the good work on your end!

OpenStudy (anonymous):

I can give it a shot. Let's take a look.